Riboflavin protein solution (75 μL and 0.66 mM) is added to a Neutral Red solution with a molarity of 0.042 mM. What is the total concentration of all Neutral Red species in solution (in mM)? The volume of Neutral Red solution is 75 μL (0.56 mM) mixed with 925 μL of a buffer solution.

Please let me know if you need any additional information!
Answer #1

75 μL of 0.56 mM Neutral Red is mixed with 925 μL buffer solution.

Total volume of the solution = (75 + 925) μL = 1000 μL.

Use the dilution equation to determine the concentration of Neutral Red.

M1V1 = M2V2

where M1 = 0.56 mM; V1 = 75 μL and V2 = 1000 μL.

Plug in values and get

(0.56 mM)*(75 μL) = M2*(1000 μL)

=====> M2 = (75 μL)*(0.56 mM)/(1000 μL)

=====> M2 = 0.042 mM.

75 μL Riboflavin is added to this solution.

Total volume of the solution becomes (1000 + 75) μL = 1075 μL.

Again, use the dilution equation.

M1V1 = M2V2

where M1 = 0.042 mM; V1 = 1000 μL and V2 = 1075 μL.

Plug in values and get

(0.042 mM)*(1000 μL) = M2*(1075 μL)

=====> M2 = (1000 μL)*(0.042 mM)/(1075 μL)

=====> M2 = 0.039 mM.

The concentration of Neutral Red in the final solution is 0.039 mM (ans).